Matrix B

Textiles (Labor-hours/unit of Output)

Wine (Labor-hours/Unit of output)

Portugal

90

80

England

100

120

(Numbers in the above matrix reflect labor hours per unit of output)

What would you recommend for the above matrix??? Reflect on all pertinent points of our analysis.
Answer #1

First, we compute Opportunity cost (OC) as follows.

In Portugal,

OC of Textile = 90/80 = 1.125 Wine

OC of Wine = 80/90 = 0.890 Textile

In England,

OC of Textile = 100/120 = 0.83 Wine

OC of Wine = 120/100 = 1.20 Textile

Since England can produce Textile at lower OC (0.83 < 1.125), England has comparative advantage in and should specialize in Textile.

Since Portugal can produce Wine at lower OC (0.890 < 1.20), Portugal has comparative advantage in and should specialize in Wine.

So, after trade, England should export Textile and import Wine, and Portugal should export Wine and import Textile.

Trade is mutually beneficial if terms of trade (TOT) lies between the OC in each country. So acceptable TOT are:

0.83 Wine < TOT of 1 unit of Textile < 1.125 Wine, or equivalently,

0.89 Textile < TOT of 1 unit of Wine < 1.20 Textile.